# NovelQA: Benchmarking Question Answering on Documents Exceeding 200K Tokens

### v2

@@ 4 Experiments
Table 1: Evaluation of Long-Context LLMs on NovelQA. This table presents the performance of four long-context LLMs, including both commercial models (GPT-4, Claude 2.1 and Claude 3 Sonnet) and open-source, locally deployed models (InternLM2-Chat-7b/20b). Accuracy percentages are reported under two testing scenarios: multichoice and generative. The Max Length column denotes the maximum token length of each model.
Max Length Multi-choice Generative
@@
Claude 3 200K 71.11 53.66
Claude 2.1 200K 66.84 46.04
@@
We focus on long-context models meeting three criteria: a context window of at least 128,000 tokens, accessibility via a full API or public release, and chat functionality. For commercial models, our selection includes GPT-4-128K (OpenAI, 2023b) and Claude 2.1-200K (Anthropic, 2023). Among open-source options, we evaluated models like InternLM2-chat (Team, 2023).
@@ 4.1 Implementations
Settings To thoroughly test the abilities of these LLMs, we employ two evaluation settings: a generative setting where models directly generate short answers, and a multi-choice setting with four provided options.
Prompts We use uniform prompts for all LLMs, with a start part, novel content, questions, choices in the multi-choice setting, and end part. The prompt structure is shown in Appendix Table 9.
@@
Evaluating Generative Results: Following the findings in Wang et al. (2023) which highlight GPT-4’s proficiency in assessing the accuracy of short machine-generated answers, we employ GPT-4 (gpt-4-0125-preview) for the evaluation of generative responses in our study, which is also applied in other long-range benchmark studies (An et al., 2023; Li et al., 2023a). We further conducted a human evaluation on 800 pieces of generative outputs and carried an inter-evaluator agreement test between two human evaluators and the GPT-4 evaluator. The result of 89.25% in Cohen’s Kappa indicates a high agreement towards the GPT-4 evaluating results (details in Appendix D.4).
Commercial LLMs: The APIs of commercial LLMs utilized are gpt-4-0125-preview [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo, Claude 2.1 and Claude 3-sonnet [^4] https://aws.amazon.com/bedrock/claude/.
Open-source LLMs: Running long-context LLMs on extremely long inputs, such as 200K tokens, is a challenge due to the immense GPU memory required, for example, it takes roughly 2.5T memory to calculate one attention matrix for a 7B model with a 200K-token input, while our local device is a 4 $\times$ 80G A100. To address this, we utilize the LMDeploy (Contributors, 2023) (based on Dynamic NTK emozilla (2023)) for memory reduction, which is only compatible with several LLMs. Therefore, we choose InternLM2-Chat-7b-200K and InternLM2-Chat-20b-200K for our experiments. Attempts to run models like Yarn-Mistral-7b-128K (Peng et al., 2023), chatglm3-6b-128K (Zeng et al., 2022), XVERSE-13B-256K (xverse, 2023) and Yi-6B-200K (01-ai, 2023) on inputs exceeding 100K tokens are unsuccessful. The challenges range from GPU memory overflow due to incompatibility with LMDeploy, to models imitating the input content, or continuing to generate following the inputs without answering the questions. However, reducing input lengths to under 50K tokens can normalize most outputs, following the given instructions. This situation underscores the current limitations in technical support for processing and generating content based on extremely long contexts.
Additional Details Given the cost of running long-context APIs, we request that each model respond to all questions for a given book in a single session. To ensure fair comparisons, local-deployed LLMs also answer all questions for a book at once. We set ‘temperature = 0’ to eliminate randomness and keep other hyper-parameters default.
Human Performance: As most of the annotators are from the same university and share common courses and projects, several books have been read by more than one student. We selected such books and have had the readers engaged in a two-round answering process on novels they had not previously annotated. The first round was in a generative setting, and the second round was in a multichoice setting. This process was conducted on 5 novels with a total of 100 questions. The result shows that human performance scored 90 in the generative setting and 97 in the multiple-choice setting.
@@ 4.2 Main Results
We present the main results in Table 12. Even the highest scores (71.80% and 46.88% for GPT-4 in generative and multi-choice settings, respectively) suggest there is considerable room for improvement in long-context understanding compared with human readers. This is especially true in the generative setting where understanding and recall over long contexts are more challenging. Additionally, commercial models (GPT-4 and Claude 3/2.1) outperform open-source models (InternLM2-Chat-7b and 20b) in this benchmark. All models show a drop in performance in the generative setting compared to the multi-choice setting. This indicates the increased challenge in generating a correct answer from scratch, as opposed to selecting from provided options.
We have also observed three typical errors, hallucination, overlooking, and miscounting, a detailed analysis is conducted in Appendix Sec E.4. In addition, we have also test the extent to which models rely on the content of novels versus using their internal knowledge to answer questions, and the details can be found in Appendix Sec E.5.
@@ 4.3 Results by the Question Type
An in-depth analysis of model performance across question types reveals nuanced insights into their comprehension abilities in both generative and multi-choice settings, detailed in Table 2 and Appendix E.2, respectively. This analysis not only highlights the models’ weaknesses across different formats but also illuminates the challenges in narrative comprehension, contributing to both NLP and computational literary research.
@@
Figure 3: Analysis of Accuracy in Generative Setting by Absolute and Relative Token Positions: The two figures above illustrates the accuracy, plotted against the absolute token indexes (left) and the percentage position (right) of each question’s relevant evidences in the novels. The x-axis of the absolute token position figure (left), reflecting token indexes, is folded on the right due to the long-tails.
4.4 Results by the Position
@@
Absolute Position Analysis: In the generative setting, as depicted in Figure 3 (left), all evaluated models show improved performance on questions where the necessary evidence is located before the 100K token mark. This trend highlights a challenge for LLMs in accessing and processing information beyond this threshold, suggesting a diminished capacity to handle very long inputs. The multi-choice setting, detailed in Appendix E.3, follows a similar pattern, reinforcing the importance of evidence position in model performance. We also present the relationship between the accuracy and the absolute position of evidence, grouping by pre-100K and post-100K, in Appendix Table 10.
Relative Position Analysis: By normalizing the evidence positions within the entire novel, we aim to understand if the proportional location of evidence influences model accuracy. This analysis, shown in Figure 3 (right) for the generative setting and in the Appendix E.3 for the multi-choice setting, indicates that models maintain relatively consistent performance across various relative positions, suggesting that long-context LLMs’ effectiveness is not significantly affected by the evidence’s relative position within the standardized length of novels.
@@ 4.5 Evidences Recall
We evaluate the ability to recall evidence, prompting the four models above to answer the questions in NovelQA again, with printing the supporting evidence simultaneously. We then prompt GPT-4 with the generated evidence alongside the annotated evidence to obtain its evaluation on the quality of retrieved evidence pieces. The evaluating matrix consists of the following three dimensions: correctness refers to whether the retrieved evidence is the same as the annotated evidence or with a similar correct meaning; relevance indicates whether the evidence is consistent with the answer; sufficiency, whether the retrieved pieces of evidence are enough to support the answer. Each dimension is scored between 0 and 100 and an average score is further obtained through calculating the algorithmic mean on these three dimensions. Prompts involved in this evaluation procedure are presented in Appendix D.5.
@@ 5 Conclusion
We introduced NovelQA, a long-range benchmark designed to assess the long-context comprehension abilities of LLMs. Utilizing representative English novels, NovelQA presents LLMs with the challenge of navigating complex, real-world texts. Our evaluations reveal that both commercial and open-source models face difficulties with detailed understanding, multi-hop reasoning, and accurately retrieving specific information from lengthy contexts, especially for lengths are 100,000. Moreover, operating LLMs on inputs exceeding 200,000 tokens faces technical challenges, notably in terms of memory requirements and associated costs.
Appendix
@@ Appendix A Limitations
Access to Close-source LLMs: One significant limitation is our inability to obtain APIs for certain close-source long-context LLMs, such as Baichuan-192K, GLM4-200K, and Moonshot-192K.
Technical Constraints for Open-source LLMs: The majority of open-source LLMs are not optimized for processing inputs exceeding 128K tokens with high GPU memory cost. This technical challenge renders direct inference on NovelQA infeasible for these models. Consequently, our experiments were limited to LLMs that have been adapted using LMDeploy.
Language Limitation: NovelQA, and all associated data are exclusively in English.
@@ Appendix B Ethics Statements
NovelQA does not contains personally identifiable information or offensive content.
Appendix C Related Work
Figure 4: Comparison between NovelQA and other long-context benchmarks highlights several key dimensions. Long dependency refers to questions that require multi-hop evidence spanning extensive text to resolve. Details pertains to questions that need answers with minimal contextual support. Reasoning involves questions without direct answers in the text, necessitating inference. NovelQA excels in these dimensions and features the longest context length among the benchmarks, catering to the current situation where a number of LLMs have a context-length equal to or over 200K.
Long-Context Language Modeling can be divided into several parts including efficient attention, preserving long-term memory like using KV cache, extrapolative positional embedding module, context pre/post-processing Huang et al. [2023], Dong et al. [2023b]. Using efficient attention can reduce computational complexity and memory usage Beltagy et al. [2020], Wang et al. [2020], Tworkowski et al. [2023], Li et al. [2023b]. Preserving KV cache or context-level cache allows models to recall and leverage past information without reprocessing, enhancing coherence over long texts Chevalier et al. [2023], Wu et al. [2022a, b], Lin et al. [2024], Zhong et al. [2023], Chen et al. [2023a], Hooper et al. [2024]. Deploying an extrapolative positional embedding to extend beyond the sequence lengths seen during training Chen et al. [2023b], Su et al. [2024]. Pre/post-processing the context can make models focus on key information Li et al. [2023c], Jiang et al. [2023], Zhang et al. [2024], Han et al. [2023]. Moreover, some training methods are put forward to support further long-context language modeling development Wang et al. [2024b], Press et al. [2021]. Since these models are more about exploring long-context language modeling without scaling in data and model size, the focus is on language modeling indicators, such as BLEU and Perplexity, NovelQA can serve as a benchmark for evaluating those long-context language modeling methods.
Appendix D Data
@@
Model Human Evaluator A Human Evaluator B Avg.
@@ D.1 Question Templates
Question templates adopted in the annotation procedure are presented in Table 5. The templates are designed using a combination of literary theories and widely-used long-context tests on LLMs. During annotation, annotators are required to fill in the missing entities (e.g., character, place, sentence) in the templates and allowed to modify the expressions to suit specific questions. This approach balances the workload of annotation and ensures a diverse set of questions.
D.2 Distribution of Input Token Count
Figure 5: Token Count Distribution in NovelQA, including Copyrighted (left) and Public Domain Only (right). The token count of both the novel and the questions are counted. The tokenization procedure is accomplished through gpt-3.5-turbo-16k tokenizer.
The input token count of each novel in NovelQA is calculated by adding the book-length to the lengths of its related questions. The distribution of the token count is illustrated in Figure 5.
D.3 Data Classification
Table 6: Distribution of Question Types in NovelQA: This table provides a breakdown of questions across different complexity categories (Multi-hop, Single-hop, Detail) and aspect categories (Time, Meaning, Span, Setting, Relation, Character, Plot).
Multi-hop Single-hop Detail Sum
Times 463 0 0 463
Meaning 34 126 206 366
Span 34 0 0 34
Setting 24 177 63 264
Relation 119 14 32 165
Character 69 255 98 422
Plot 64 414 113 591
Sum 807 986 512 2305
@@ D.3 Data Classification
D.4 Inter-Annotator Agreement Tests
We have conducted two separate Inter-Annotator Agreement (IAA) tests in this work.
The first IAA experiment focuses on evaluating the quality of annotated question-answer pairs. Annotators are required to choose books they are familiar with but have not annotated themselves to answer questions on. As annotators are mostly from the same local university and share similar courses and projects, many can find books they have read in common. We select books read by at least two annotators and have the other reader answer multichoice questions. As the respondents are quite familiar with the target novels and have a strong academic background in English Literature, it takes them around 2-3 hours or less to complete each novel.
The second IAA test assesses model answer evaluation by calculating the correlation between human evaluators and GPT-4 evaluators. Annotators also serve as human evaluators for model-generated answers to the novels they were familiar with. Their evaluations were compared to those of GPT-4 evaluators, with Cohen’s kappa score calculated to measure agreement. Human evaluators, familiar with the content, typically completed their reviews in under half an hour per novel. We also perform an inter-evaluator agreement test between human-evaluators and the GPT-4-evaluator. The results of Cohen’s Kappa score are presented in Table 4.
D.5 Prompts
@@
Appendix E Experiments and Analysis
Multichoice Generative
pre-100K post-100K pre-100K post-100K
GPT-4 83.17 59.22 53.64 32.01
Claude2.1 77.65 54.46 61.96 30.57
InternLM-7b 51.78 38.17 37.71 18.23
InternLM-20b 55.59 40.87 37.74 21.10
Weighted Avg. 66.24 48.08 46.84 25.36

### v3

@@ 4 Experiments
Table 2: Evaluation of Long-Context LLMs on NovelQA. This table presents the performance of four long-context LLMs, including both commercial models (GPT-4, GPT-4o-mini, Claude-2.1, Claude-3-Sonnet and Claude-3.5-Sonnet) and open-source, locally deployed models (InternLM2-Chat-7b/20b and Llama-3.1-8b/70b). Accuracy percentages are reported under two testing scenarios: multichoice and generative. The Max Length column denotes the maximum token length of each model.
Max Length Multichoice Generative
@@
GPT-4o-mini 128K 71.85 53.32
Claude-2.1 200K 66.84 46.04
Claude-3 200K 71.11 53.66
Claude-3.5-Sonnet 200K 77.92 62.30
@@
We focus on long-context models meeting three criteria: a context window of at least 128,000 tokens, accessibility via a full API or public release, and chat functionality. For commercial models, our selection includes GPT-4-128K (OpenAI, 2023a) and Claude 2.1-200K (Anthropic, 2023). Among open-source options, we evaluated models like InternLM2-chat (Team, 2023).
@@ 4.1 Implementations
Settings We employ two evaluation settings: a generative setting where models directly generate short answers, and a multichoice setting with four provided options.
Prompts We use uniform prompts for all LLMs, with a start part, novel content, questions, choices in the multichoice setting, and an end part. The prompt structure is shown in Appendix Table 11.
@@
Evaluating Generative Results Previous researches have proved the ability of LLMs in evaluating the machine-generated answers align with human judgements (Wang et al., 2023; An et al., 2023; Li et al., 2023a) After a pilot study showing that the models show no several preference towards the answers generated by its own kinds (See Appendix C.4.2, we choose GPT-4 to evaluate our generated results. We further conducted a human evaluation on 800 pieces of generative outputs and carried out an inter-evaluator agreement (IEA) test between two human evaluators and the GPT-4 evaluator. In the IEA test, annotators also serve as human evaluators of the novels they were familiar with. Their evaluations were compared to those of GPT-4 evaluators, with Cohen’s kappa score calculated to measure agreement. As shown in Table 7, the result of 89.25% in Cohen’s Kappa indicates a high agreement towards the GPT-4 evaluating results. NovelQA primarily consists of objective questions, which have clear, verifiable answers with a factual basis in the text. This objectivity significantly reduces the impact of model bias in LLM-as-Judge evaluation
Commercial LLMs The APIs of commercial LLMs utilized are gpt-4-0125-preview, gpt-4o-mini, Claude-2.1, Claude-3-sonnet and Claude-3.5-sonnet.
Open-source LLMs Running long-context LLMs on extremely long inputs, such as 200K tokens, is a challenge due to the immense GPU memory required, for example, it takes roughly 2.5T memory to calculate one attention matrix for a 7B model with a 200K-token input, while our local device is a 4 $\times$ 80G A100. To address this, we utilize the LMDeploy (Contributors, 2023) (based on Dynamic NTK (emozilla, 2023)) and vLLM (Kwon et al., 2023) for memory and time reduction, which is only compatible with several LLMs. Therefore, we choose InternLM2-Chat-7b-200K, InternLM2-Chat-20b-200K, Llama-3.1-8b, and Llama-3.1-70b for our experiments.
Human Performance We selected books that have been read by multiple annotators, and then have had the readers engage in a two-round answering process on novels they had not previously annotated. The first round was in a generative setting, and the second round was in a multichoice setting. This process was conducted on 5 novels with a total of 100 questions. The result shows that human performance scored 90 in the generative setting and 97 in the multiple-choice setting.
@@ 4.2 Main Results
We present the main results in Table 2. Even the highest scores (71.80% and 46.88% for GPT-4 in generative and multichoice settings, respectively) suggest there is considerable room for improvement in long-context understanding compared with human readers. This is especially true in the generative setting where understanding and recall over long contexts are more challenging. Additionally, commercial models (Claude-3.5-Sonnet and GPT-4o-mini) outperform open-source models in this benchmark. Among closed-source models, LLaMA family models outperform the InternLM models. All models show a drop in performance in the generative setting compared to the multichoice setting. This indicates the increased challenge in generating a correct answer from scratch, as opposed to selecting from provided options.
We appended a typical error analysis in Appendix Sec C.4.6.
@@ 4.3 Results by the Question Type
An in-depth analysis of model performance across question types reveals nuanced insights into their comprehension abilities in both generative and multichoice settings, detailed in Table 3 and Appendix C.4.3, respectively. This analysis not only highlights the models’ weaknesses across different formats but also illuminates the challenges in narrative comprehension, contributing to both NLP and computational literary research.
@@
Figure 4: Analysis of Accuracy in Generative Setting by Absolute and Relative Token Positions: The two figures above illustrate the accuracy, plotted against the absolute token indexes (left) and the percentage position (right) of each question’s relevant evidences. The x-axis of the absolute token position figure (left), reflecting token indexes, is folded on the right due to the long-tails.
4.4 Results by Position
Multichoice Generative
pre-100K post-100K pre-100K post-100K
GPT-4 83.17 59.22 53.64 32.01
Claude2.1 77.65 54.46 61.96 30.57
InternLM-7b 51.78 38.17 37.71 18.23
InternLM-20b 55.59 40.87 37.74 21.10
Weighted Avg. 66.24 48.08 46.84 25.36
@@ 4.4 Results by Position
Absolute Position In the generative setting, as depicted in Figure 4 (left), all evaluated models show improved performance on questions where the necessary evidence is located before the 100K token mark. This trend highlights a challenge for LLMs in accessing and processing information beyond this threshold, suggesting a diminished capacity to handle very long inputs. The multichoice setting, detailed in Appendix C.4.4, follows a similar pattern, reinforcing the importance of evidence position in model performance. We also present the relationship between the accuracy and the absolute position of evidence, grouping by pre-100K and post-100K, in Table 4.
Relative Position By normalizing the evidence positions within the entire novel, we aim to understand if the proportional location of evidence influences model accuracy. This analysis, shown in Figure 4 (right) for the generative setting and in the Appendix C.4.4 for the multichoice setting, indicates that models maintain relatively consistent performance across various relative positions, suggesting that long-context LLMs’ effectiveness is not significantly affected by the evidence’s relative position within the standardized length of novels.
@@
This analysis highlights the critical role of absolute evidence positioning in determining the accuracy of LLMs in processing long texts. Challenges arise when context beyond a specific token threshold. Conversely, the relative position within a normalized length has a minimal effect.
@@ 4.5 Evidences Recall
We evaluate the ability to recall evidence, prompting the four models above to answer the questions in NovelQA again, with printing the supporting evidence simultaneously. We then prompt GPT-4 with the generated evidence alongside the annotated evidence to obtain its evaluation on the quality of retrieved evidence pieces. The evaluating matrix consists of the following three dimensions: correctness refers to whether the retrieved evidence is the same as the annotated evidence or with a similar correct meaning; relevance indicates whether the evidence is consistent with the answer; sufficiency, whether the retrieved pieces of evidence are enough to support the answer. Each dimension is scored between 0 and 100 and an average score is further obtained through calculating the algorithmic mean on these three dimensions. Prompts involved in this evaluation procedure are presented in Appendix C.2.3.
@@
Close-book QA
Multichoice Generative
GPT-4 60.94 (-9.06) 34.30 (-13.58)
Claude 2.1 51.77 (-15.01) 22.36 (-20.68)
InternLM2-Chat-7b 33.58 (-10.71) 14.12 (-16.81)
InternLM2-Chat-20b 33.05 (-16.13) 15.51 (-16.86)
4.6 Close-book Question Answering
We employ a Close-book QA scenario to assess the extent to which models rely on the content of novels versus using their internal knowledge to answer questions. In this approach, the models are not given access to the text of the novels and must rely solely on their pre-existing knowledge to provide answers. Given that our selected novels are well-known and representative of their genres, it’s inevitable that LLMs have encountered their texts during training and retained some of their content. The results of this evaluation are presented in Table 6. Models like GPT-4 and Claude 2.1 achieve notable scores in the Close-book setting (60.94% and 51.77% in multichoice, 34.30% and 22.36% in generative, respectively) indicating that they have internalized significant portions of the novels’ content during training. This internal knowledge allows them to perform reasonably well even without direct access to the text. The difference in performance between the Close-book and standard settings underscores the challenges in long-context understanding. While models can retain and recall information from well-known texts, their ability to comprehend and use such information to answer questions accurately diminishes in the absence of the text. This suggests that long-context understanding, as measured by the main results, might still be more challenging than it appears, as models benefit from having the text directly available.
@@ 5 Conclusion
We introduced NovelQA, a long-range benchmark designed to assess the long-context comprehension abilities of LLMs. Utilizing representative English novels, NovelQA presents LLMs with the challenge of navigating complex, real-world texts. Our evaluations reveal that both commercial and open-source models face challenges with detailed understanding, multi-hop reasoning, and accurately retrieving specific information from lengthy contexts, especially for lengths are 100,000. Moreover, operating LLMs on inputs exceeding 200,000 tokens faces technical challenges, notably in terms of memory requirements and associated costs.
Acknowledgment
We appreciate the contributions of our annotators: Jingyi Wu, Ziyue Li, Qiangyi Zhu, Xinyao Feng, Zhangrui Xu, Xiao Wu, Yu Qiu, Hanxin Ye, Mengting Wei, Hanyu Cai, and Yaokai Tu from Hangzhou Normal University; Ziyang Zhu, and Xinyu Zhu from Ningbo University; and Rizwan Abbas, Khadija Azhar, and Snawar Hussainfrom Zhejiang University. This publication has emanated from research conducted with the financial support of the National Natural Science Foundation of China (NSFC No.62161160339)
@@ Appendix A Limitations
Genre Limitation: NovelQA focus on the novel genre. The contents of several popular novels might be mentioned in existing datasets, resulting in the risk of containing the correct answers of our copyright protected books, although this dataset has managed to contain non-trivial details and reasoning questions to avoid this risk.
Language Limitation: NovelQA, and all associated data are exclusively in English. The following researches may consider to extend the language covered.
Truncation: Truncation is a standard practice for adapting longer texts to fit within the constrained context windows of LLMs, especially when evaluating their performance on long-context tasks. While it’s possible that some evidence may fall within the truncated parts.
@@ Appendix B Ethics Statements
NovelQA does not contain any personally identifiable information or offensive content.
Appendix C Appendix
C.1 Related Works
Long-Context Language Modeling can be divided into several parts including efficient attention, preserving long-term memory like using KV cache, extrapolative positional embedding module, context pre/post-processing (Huang et al., 2023; Dong et al., 2023b). Using efficient attention can reduce computational complexity and memory usage (Beltagy et al., 2020; Wang et al., 2020; Tworkowski et al., 2023; Li et al., 2023b). Preserving KV cache or context-level cache allows models to recall and leverage past information without reprocessing, enhancing coherence over long texts (Chevalier et al., 2023; Wu et al., 2022a; b; Lin et al., 2024; Zhong et al., 2023; Chen et al., 2023a; Hooper et al., 2024). Deploying an extrapolative positional embedding can extend beyond the sequence lengths seen during training (Chen et al., 2023b; Su et al., 2024). Pre/post-processing the context can make models focus on key information (Li et al., 2023c; Jiang et al., 2023; Zhang et al., 2024; Han et al., 2023). Moreover, some training methods are put forward to support further long-context language modeling development (Wang et al., 2024a; Press et al., 2021). Since these models are more about exploring long-context language modeling without scaling in data and model size, the focus is on language modeling indicators, such as BLEU and Perplexity, NovelQA can serve as a benchmark for evaluating those long-context language modeling methods.
Figure 5: Comparison between NovelQA and other long-context benchmarks highlights several key dimensions. Long dependency refers to questions that require multi-hop evidence spanning extensive text to resolve. Details pertains to questions that need answers with minimal contextual support. Reasoning involves questions without direct answers in the text, necessitating inference. NovelQA excels in these dimensions and features the longest context length among the benchmarks, catering to the current situation where a number of LLMs have a context length equal to or over 200K.
C.2 Data
@@
Model To Human Evaluator A To Human Evaluator B Avg.
@@ C.2.1 Question Templates
Question templates adopted in the annotation procedure are presented in Table 8. The templates are designed using a combination of literary theories and widely-used long-context tests on LLMs. During annotation, annotators are required to fill in the missing entities (e.g., character, place, sentence) in the templates and are allowed to modify the expressions to suit specific questions. This approach balances the workload of annotation and ensures a diverse set of questions.
C.2.2 Data Classification
@@
C.2.3 Prompts
@@
C.2.4 The Length Distribution of Evidences
Range 0~20K 20~40K 40~60K 60~80K 80~100K 100~130K 130~170K 170~210K >210K
Count 959 631 446 359 249 239 214 209 226
Percentage 27.15 17.87 12.63 10.16 7.05 6.77 6.06 5.91 6.40
Table 12: The distribution of evidences by absolute token index.
Range 0~10 20 30 40 50 60 70 80 90 100
Count 889 431 360 326 288 271 274 250 205 238
Percentage 25.17 12.20 10.19 9.23 8.15 7.67 7.76 7.08 5.80 6.74
Table 13: The distribution of evidences by relative token index.
Of the 3,666 evidence instances, 134 are summaries or comments without exact matches in the document. The distribution of the remaining 3,532 instances is as follows: Absolute distribution (by absolute token index) can be found in Table 12: Approximately 18.4% of the evidence is located beyond the 130K token mark (outside GPT-4’s context window), and 6.4% is beyond 210K tokens (outside Claude and InternLM’s context windows).
Given the varying lengths of novels, we also analyzed the relative position distribution (by relative token position: token position / total tokens in the corresponding book) in Table 13:
Our analysis reveals that the highest proportion of evidence instances (25.17%) occurs within the first 10% of the novels, while the distribution across the middle sections is relatively uniform. This concentration in the early parts of the novels can be attributed to the initial introduction of characters and plot elements. The first occurrence of evidence related to these introductions naturally falls in the earlier sections of the novels. It’s important to note that during the question formulation process, we did not deliberately adjust the distribution of questions. The observed pattern in answer locations emerges naturally from the narrative structure of the novels.
C.3 Multichoice Distractors Generation
Our pilot study compared distractors generated by both GPT-4 and Claude 2.1. Interestingly, GPT-4 generated slightly more challenging distractors - both models scored approximately only 0.5% lower on GPT-4’s distractors compared to Claude 2.1’s. Thus, we consider the bias caused by the model choice not significant in the distractors generation, and safely chose the choices generated by GPT-4.
C.4 Experiments and Analysis